# A framework for fluctuating times and counting observables in stochastic excursions

### v1

@@ Abstract
Many natural systems exhibit dynamics characterized by alternating phases. Describing the fluctuations of such systems over stochastic trajectories is necessary across diverse fields, from biological motors to quantum thermal machines. In an accompanying Letter, we introduced the notion of stochastic excursions—a framework to analyze far from equilibrium processes via sub-trajectories. Through counting observables, this framework captures finite-time fluctuations and trajectory-level behavior, which provides insights into thermodynamical trade-offs between thermodynamic quantities of interest, such as entropy production and dynamical activity. In this work, we enhance this formalism by providing a suite of technical results on how to efficiently compute excursion-related quantities. Our analytical results provide explicit formulas for general moments of counting variables and excursion duration, as well as their covariance and conditional moments. We show that excursion statistics recover full counting statistics results, and uncover a relation between fluctuations of counting observables at single-excursion level and the steady state diffusion coefficient (noise). We also discuss a fluctuation theorem for individual excursions. In addition, we explore how analyzing excursions and using the results developed here can yield insights into three problems of interest: the three-qubit absorption refrigerator, cellular sensing, and birth-and-death processes.
@@ I Introduction
A signature of many natural phenomena is the alternating behavior between an “inactive” phase $A$ and an “active” phase $B$. These kinds of systems transition from $A$ to $B$ when activated (e.g., by an external energy input), perform some arbitrary task in $B$, and eventually return to $A$. This type of trajectory is characterized by the concept of stochastic excursions, which is a framework for analyzing sub-trajectories of processes far from equilibrium Fiusa2025 . Problems of this form appear across diverse fields, including biology bergPhysicsChemoreception1977 ; Mora2010 ; Mora2010 ; Harvey2023 ; Fancher2020 ; Barato2015 ; Lang2014 , physics Prech2024 ; Fiusa2024 ; Hegde2025 , chemistry bialekPhysicalLimitsBiochemical2005 , and applied mathematics kleinrock1974queueing ; Lindley1952 ; gross2011fundamentals ; cohen1969single ; Bhat2008 . While the concept of brownian excursions has been explored in other contexts Chung1976 ; Majumdar2005 ; Majumdar2008 ; Majumdar2015 ; Takacs1991 ; Mazzolo2017 , our work introduces a special type of filtering of trajectories which can be thought of as a generalization. Furthermore, most theoretical frameworks are model-based and have focused on questions related to first-passage times Yao1985 ; Rubino1989 ; Kook1993 ; Takacs1974 ; Garrahan2017 ; Bebon2024 . Those problems can be recast as questions related to excursion times.
@@
This work is a longer companion to the Letter Fiusa2025 , where we introduced the formalism, showcased its ubiquity across numerous problems of interest, and discussed the main results of this framework. Our goal in this accompanying paper is to construct the formalism and provide the tools and technical results to efficiently compute excursion-related quantities. We also expand on applications by showing how we use the formalism developed herein to characterize fluctuations in the three-qubit absorption refrigerator, a cellular sensing problem, and a typical birth-and-death process.
@@
The organization of the present work is as follows. In Sec. II, we introduce the formalism of stochastic excursions and counting observables. In Sec. III, we explore the statistics of counting observables and present our main technical results, including analytical formulas for the moments of counting observables, excursion durations, and their correlations. Furthermore, we use the results derived in the section to connect our framework to full counting statistics landi2024a ; Campisi2011 ; Esposito2007 ; Esposito2009 ; Brandes2008 ; Saito2008 , providing formulas for average currents and diffusion coefficients. In Sec. IV, we derive explicit results for excursion activity and specific transition counts. In Sec. V, we show a fluctuation theorem Jarzynski1997 ; Crooks1999 ; Talkner2007a ; Talkner2007b ; Talkner2008 ; Hasegawa2019a ; Hasegawa2019b ; Campisi2009 ; Campisi2010 ; Campisi2014 ; Andrieux2009 ; Jarzynski2004 ; Jarzynski2011 ; Seifert2012 ; Esposito2009 ; Harris2007 ; Horowitz2019 ; Nakamura2010 for individual excursions and relate it to existing results in the literature. In Sec. VI, we apply our framework to three problems of interest: the three-qubit absorption refrigerator, cellular sensing, and birth-and-death processes. Finally, in Sec. VII, we discuss the broad implications of our findings and outline future directions.
@@ II.1 Regions $A$ and $B$
Figure 1: A system undergoes stochastic dynamics on a finite set of states. The states are split into two regions, called $A$ (grey) and $B$ (green). An excursion starts with a transition from $A\to B$ and ends with the first transition back from $B\to A$ (red arrows). The states of region $A$ can be spread out across the space state.
Suppose we split the alphabet of states in two arbitrary regions, denoted by $A$ and $B$, see Fig 1. The transition matrix is thus split into blocks, as
@@
as well as the vector $\langle 1|=(\langle 1_{A}|\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \langle 1_{B}|)$. This provides the following identities
@@
$\langle 1_{A}|\mathbb{W}_{\!A}$ $=-\langle 1_{B}|{W}_{BA},\qquad\leavevmode\nobreak\ \leavevmode\nobreak\ \langle 1_{A}|{W}_{AB}=-\langle 1_{B}|\mathbb{W}_{\!B}.$
@@ II.2 Excursions
$P_{\tt exc}=CW_{\ell_{N}y_{A}z_{N}}e^{-\Gamma_{z_{N}}\tau_{N}}W_{\ell_{N-1}z_{N}z_{N-1}}\ldots e^{-\Gamma_{z_{1}}\tau_{1}}W_{\ell_{0}z_{1}x_{A}},$ (7)
where $z_{i}$ are all states within $B$ and $C$ is a normalization constant (see below). The factors of $e^{-\Gamma_{x}\tau}$ represent the residence times in state $x$ before jumping, while $W_{\ell x^{\prime}x}$ represent the transition rates between the states $x$ and $x^{\prime}$ due to $\ell$. The number of states $z_{1},z_{2},\ldots,z_{N}$ that the system visits in $B$ is arbitrary but must be at least one, so the smallest possible excursion will have the form $x_{A}\to z_{1}\to y_{A}$.
Let $P[x_{A}\to z_{B}]$ denote the probability that an excursion starts with a jump $x_{A}\to z_{B}$ given that the it starts in $x_{A}$:
@@
where $\bra{1_{B}}$ | is a vector with all entries equal to 1. The denominator $\mathcal{K}_{BA}^{x_{A}}:=\braket{1_{B}}{W_{BA}}{x_{A}}$ ensures that the probability is properly normalized. We see that it corresponds to the dynamical activity from state $x_{A}$ to region $B$; more specifically, it is the average number of transitions per unit time from $x_{A}$ to $B$, and represents how frequently excursions occur stemming from $x_{A}$.
@@
$P[y_{A}\mid z_{B}]=\braket{y_{A}}{W_{AB}M}{z_{B}},$ (9)
where $M:=-\mathbb{W}_{B}^{-1}$ accounts for the variable time that the system spends in region $B$ [c.f. Eq. (14)]. Multiplying Eqs. (8) and (9) and summing over all states |$\ket{z_{B}}$ yields the probability that an excursion ends in $y_{A}$ given that it started in $x_{A}$,
$P[x_{A}\to y_{A}|x_{A}]=\frac{\braket{y_{A}}{W_{AB}MW_{BA}}{x_{A}}}{\mathcal{K}_{BA}^{x_{A}}}.$ (10)
@@
$C_{x_{A}\to y_{A}}^{-1}=P[x_{A}\to y_{A}]\mathcal{K}_{BA}^{x_{A}}.$ (11)
We therefore see that $C_{x_{A}\to y_{A}}^{-1}$ is related to the probability of the excursion ending in $y_{A}$ conditioned on its start in $x_{A}$, and the dynamical term $\mathcal{K}_{BA}^{x_{A}}$. This attributes physical meaning to the normalization constant $C$ that appears in several calculations throughout this text. When it is clear from the context, we sometimes omit the labels “$x_{A}\to y_{A}$” to simplify the notation.
@@
The duration of an excursion is a random variable, which we denote by $\hat{\mathcal{T}}$. This is a first-passage time. Its PDF can be obtained directly from Eq. (7) as
$P\big{(}\hat{\mathcal{T}}=t\big{)}=E\Biggl{[}\delta\Bigg{(}t-\sum_{j=1}^{N}\tau_{j}\Bigg{)}\Biggr{]}.$ (13)
@@
$P\big{(}\hat{\mathcal{T}}=t\big{)}=\leavevmode\nobreak\ C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB}e^{\mathbb{W}_{\!B}t}W_{BA}|x_{A}\rangle.$ (14)
This is a minor generalization of known results on first passage times skinner2021 .
@@ II.3 Counting observables
For a given single excursion, we define for each pair $(x,y)$ a random variable $\hat{N}_{\ell xy}$ that counts how many times the transition $y\to_{\ell}x$ was observed (with $x,y\in A\cup B$). A counting observable is defined as a linear combination of these counting variables:
@@
with generic weights $\nu_{\ell xy}$ landi2024a . Counting observables are defined irrespective of which state the excursion began and ended. In standard full counting statistics (FCS), a counting observable $\hat{Q}(t)$ is defined so as to count the transitions over a fixed interval $[0,t]$. The connection between our results and FCS is discussed in Sec. III.6. Here, the new element we introduce is the concept of a counting observable $\hat{Q}$ that counts events during a single excursion, whose duration is also random.
The simplest example of a counting observable is counting how many transitions occur in total. This is achieved by setting $\nu_{\ell xy}=1$ in Eq. (15). This counting observable is referred as the excursion activity $\hat{\mathcal{A}}$, it is related to the more usual dynamical activity (freneticity) which characterizes the number of transitions per unit time. The difference is that instead of characterizing activity per unit time, excursion activity accounts for activity per excursion. We provide analytical formulas for the excursion activity in Sec. IV.1.
With different choices for the $\nu_{\ell xy}$ we can assess a broad variety of physical quantities. A typical quantity of interest is the number of times a specific transition $z^{\prime}\to z$ inside region $B$ occurs. This is obtained by setting $\nu_{\ell zz^{\prime}}=1$ and all other weights as zero. We provide analytical formulas for this case in Sec. IV.2. Similarly, $\nu_{\ell xy}=1\,\forall x,y$ for some $\ell$, can address the number of transitions of a given type. This can evaluate, for example, the activity of some hot reservoir in contrast to some other cold reservoir.
@@
where $E_{x/y}$ are the energies of levels $x/y$ and $\beta_{\ell}$ is the inverse temperature of the reservoir associated with transition $\ell$. The counting observable representing the entropy produced in the process can then be built with weights $\nu_{\ell xy}=-\beta_{\ell}(E_{x}-E_{y})$, which results in
@@
Other thermodynamic quantities can be constructed similarly. In Sec. V, we show that the entropy production satisfies a fluctuation theorem at the excursion level, which recovers known results in trajectory-level fluctuations seifertEntropyProductionStochastic2005a .
@@ III Statistics of counting observables within an excursion
$(\mathbb{W}_{\bm{\xi}})_{xy}=\begin{cases}\sum_{\ell}W_{\ell xy}e^{i\sum_{\alpha}\nu_{\ell xy}^{\alpha}\xi_{\alpha}}&x\neq y\\[5.69046pt] -\Gamma_{x}&x=y,\end{cases}$ (18)
with $r$ counting fields $\xi_{\alpha}$. For an excursion starting in $x_{A}$ and ending in $y_{A}$, the joint probability that each observable $\hat{Q}_{\alpha}$ takes on a value $q_{\alpha}$ and that total excursion time is $\hat{\mathcal{T}}=t$ reads
$P(\bm{q},t)=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi_{1}\ldots d\xi_{r}}{(2\pi)^{r}}\langle y_{A}|W_{AB\bm{\xi}}e^{\mathbb{W}_{\!B\bm{\xi}}t}W_{BA\bm{\xi}}|x_{A}\rangle e^{-i\bm{q}\cdot\bm{\xi}},$ (19)
@@
$P(\bm{q})=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi_{1}\ldots d\xi_{r}}{(2\pi)^{r}}\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}W_{BA\bm{\xi}}|x_{A}\rangle e^{-i\bm{q}\cdot\bm{\xi}},$ (20)
@@
The results stated in this section so far rely on a specific choice of initial and final states, $x_{A}$ and $y_{A}$, within region $A$. Very often, one is interested only in the statistics of excursions, irrespective of where exactly it starts or ends. This is true, for example, when we want to analyze the long-time statistics over many excursions. In this case, results from Eqs. (14), (19), and (20) remain valid provided that initial states $x_{A}$ are sampled from the the steady state distribution $|p_{A}^{\rm ss}\rangle$ [Eq. (5)], and we sum over all possibilities for the final state, which amounts to replacing $\langle y_{A}|$ by $\langle 1_{A}|$, which is the vector of length $|A|$ with all entries equal to 1.

### v2

@@ Abstract
Many natural systems exhibit dynamics characterized by alternating phases or recurring sets of states. Describing the fluctuations of such systems over stochastic trajectories is necessary across diverse fields, from biological motors to quantum thermal machines. In an accompanying Letter, we introduced the notion of stochastic excursions—a framework to analyze out of equilibrium processes via sub-trajectories. Through counting observables, this framework captures finite-time fluctuations and trajectory-level behavior, which provides insights into thermodynamical trade-offs between thermodynamic quantities of interest, such as entropy production and dynamical activity. In this work, we enhance this formalism by providing a suite of technical results on how to efficiently compute excursion-related quantities. Our analytical results provide explicit formulas for general moments of counting variables and excursion duration, as well as their covariance and conditional moments. We show that excursion statistics recover full counting statistics results, and uncover a relation between fluctuations of counting observables at single-excursion level and the steady state diffusion coefficient (noise). We also discuss a fluctuation theorem for individual excursions and show that it implies a modified thermodynamic uncertainty relation at the excursion level. In addition, we explore how analyzing excursions and using the results developed here can yield insights into three problems of interest: the three-qubit absorption refrigerator, cellular sensing, and birth-and-death processes.
@@ I Introduction
A signature of many natural phenomena is the alternating behavior between an “inactive” phase $A$ and an “active” phase $B$. These kinds of systems transition from $A$ to $B$ when activated (e.g., by an external energy input), perform some arbitrary task in $B$, and eventually return to $A$. This type of trajectory is characterized by the concept of stochastic excursions, which is a framework for analyzing sub-trajectories of out of equilibrium Markov jump processes Fiusa2025 . Systems with a recurring set of states are naturally phrased in the stochastic excursion language. Problems of this form appear across diverse fields, including biology bergPhysicsChemoreception1977 ; Mora2010 ; Mora2010 ; Harvey2023 ; Fancher2020 ; Barato2015 ; Lang2014 ; Barato2016 ; Barato2014 , molecular machines Seifert2012 ; Kumar2010 ; Kolomeisky2007 ; Murugan2014 , heat engines and physics Brandner2015 ; Pietzonka2018 ; Prech2024 ; Fiusa2024 ; Hegde2025 , chemistry bialekPhysicalLimitsBiochemical2005 , and applied mathematics kleinrock1974queueing ; Lindley1952 ; gross2011fundamentals ; cohen1969single ; Bhat2008 . Furthermore, most theoretical frameworks proposed to study such systems are model-based and have focused on questions related to first-passage times Yao1985 ; Rubino1989 ; Kook1993 ; Takacs1974 ; Garrahan2017 ; Bebon2024 . Those problems can be recast as questions related to excursion times. While the concept of excursions has been explored in other contexts Chung1976 ; Majumdar2005 ; Majumdar2008 ; Majumdar2015 ; Takacs1991 ; Mazzolo2017 ; Zapperi2005 ; Papanikolaou2011 ; Baldassarri2021 , our work is unrelated because it introduces a special type of filtering of trajectories and does not address Brownian/Langevin type-dynamics, but rather Markov jump processes.
@@
This work is a longer companion to the Letter Fiusa2025 , where we introduced the formalism, showcased its ubiquity across numerous problems of interest, and discussed the main results of this framework. Our goal in this accompanying paper is to construct the formalism and provide the tools and technical results to efficiently compute excursion-related quantities. At the core of our framework lies the concept of counting observables and their statistics during excursions. These observables are constructed by counting the number of times a specific transition took place within an excursion, and appending the appropriate weights. A wide range of observables such as dynamical activity and thermodynamic quantities, such as heat, work, currents, entropy production, are constructed this way. In this longer companion paper, we provide a toolkit of simple formulas to evaluate the statistics of excursion durations, counting observables, and their interplay. All formulas are written as functions of elements of the stochastic transition matrix. We also show that a fluctuation theorem specialized for excursions naturally appear from the probability distribution of counting observables and local detail balance. We use this fluctuation theorem to derive a thermodynamic uncertainty relation at the excursion level. We also expand on applications by showing how we use the formalism developed herein to characterize fluctuations in the three-qubit absorption refrigerator, a cellular sensing problem, and a birth-and-death process.
@@
The organization of the present work is as follows. In Sec. II, we introduce the formalism of stochastic excursions and counting observables. In Sec. III, we explore the statistics of counting observables and present our main technical results, including analytical formulas for the moments of counting observables, excursion durations, and their correlations. Furthermore, we use the results derived in the section to connect our framework to full counting statistics landi2024a ; Campisi2011 ; Esposito2007 ; Esposito2009 ; Brandes2008 ; Saito2008 , providing formulas for average currents and diffusion coefficients. In Sec. IV, we derive explicit results for excursion activity and specific transition counts. In Sec. V, we show a fluctuation theorem Jarzynski1997 ; Crooks1999 ; Talkner2007a ; Talkner2007b ; Talkner2008 ; Hasegawa2019a ; Hasegawa2019b ; Campisi2009 ; Campisi2010 ; Campisi2014 ; Andrieux2009 ; Jarzynski2004 ; Jarzynski2011 ; Seifert2012 ; Esposito2009 ; Harris2007 ; Horowitz2019 ; Nakamura2010 for individual excursions and use it to derive a thermodynamic uncertainty relation Gingrich2016 ; Hasegawa2019 ; Potts2019 where the fluctuations of observables along excursions in forward and backward trajectories are always tied via entropy production. In Sec. VI, we apply our framework to three problems of interest: the three-qubit absorption refrigerator, cellular sensing, and birth-and-death processes. Finally, in Sec. VII, we discuss the broad implications of our findings and outline future directions.
@@ II.1 Regions $A$ and $B$
Figure 1: A system undergoes stochastic dynamics on a finite set of states. The states are split into two regions, called AA (grey) and $B$ (green). An excursion starts with a transition from $A\to B$ and ends with the first transition back from $B\to A$ (red arrows). The states of region AA can be spread out across the space state.
Suppose we split the alphabet of states in two arbitrary regions, denoted by $A$ and $B$, see Fig 1. The transition matrix is then split into blocks, as
@@
as well as the vector $\langle 1|=(\langle 1_{A}|\penalty 10000\ \penalty 10000\ \penalty 10000\ \langle 1_{B}|)$. This provides the following identities
@@
$\langle 1_{A}|\mathbb{W}_{\!A}$ $=-\langle 1_{B}|{W}_{BA},\qquad\penalty 10000\ \penalty 10000\ \langle 1_{A}|{W}_{AB}=-\langle 1_{B}|\mathbb{W}_{\!B},$
@@ II.2 Excursions
$P_{\tt exc}=C_{x_{A}\to y_{A}}W_{\ell_{N}y_{A}z_{N}}e^{-\Gamma_{z_{N}}\tau_{N}}W_{\ell_{N-1}z_{N}z_{N-1}}\ldots e^{-\Gamma_{z_{1}}\tau_{1}}W_{\ell_{0}z_{1}x_{A}},$ (7)
where $z_{i}$ are all states within $B$ and $C_{x_{A}\to y_{A}}$ is a normalization constant that in general depends on the states that the excursion begins and ends (see below)  [^1] It is also possible to compute Eq. (7) without conditioning on any particular initial or final state, it amounts to summing (marginalizing) over all states $x_{A}$ and $y_{A}$.. The factors of $e^{-\Gamma_{x}\tau}$ represent the residence times in state $x$ before jumping, while $W_{\ell x^{\prime}x}$ represent the transition rates between the states $x$ and $x^{\prime}$ due to $\ell$. The number of states $z_{1},z_{2},\ldots,z_{N}$ that the system visits in $B$ is random but must be at least one, so the smallest possible excursion will have the form $x_{A}\to z_{1}\to y_{A}$.
Let $P[x_{A}\to z_{B}|x_{A}]$ denote the probability that an excursion starts with a jump $x_{A}\to z_{B}$ given that the it starts in $x_{A}$:
$P[x_{A}\to z_{B}|x_{A}]=\frac{\braket{z_{B}|W_{BA}|x_{A}}}{\braket{1_{B}|W_{BA}|x_{A}}},$ (8)
@@
$P[y_{A}\mid z_{B}]=\braket{y_{A}|W_{AB}M|z_{B}},$ (9)
[…]
$P[x_{A}\to y_{A}|x_{A}]=\frac{\braket{y_{A}|W_{AB}MW_{BA}|x_{A}}}{\mathcal{K}_{BA}^{x_{A}}}.$ (10)
@@
$C_{x_{A}\to y_{A}}^{-1}=P[x_{A}\to y_{A}|x_{A}]\mathcal{K}_{BA}^{x_{A}}.$ (11)
We therefore see that $C_{x_{A}\to y_{A}}^{-1}$ is related to the probability of the excursion ending in $y_{A}$ conditioned on its start in $x_{A}$, and the dynamical term $\mathcal{K}_{BA}^{x_{A}}$. This attributes physical meaning to the normalization constant $C_{x_{A}\to y_{A}}$ that was first seen in Eq. (7) and appears in several calculations throughout this text.
@@
The duration of an excursion is a random variable, which we denote by $\hat{T}$. This is a first-passage time. Its PDF can be obtained directly from Eq. (7) as
$P\big(\hat{T}=t\big)=E\Biggl[\delta\Bigg(t-\sum_{j=1}^{\hat{N}}\tau_{j}\Bigg)\Biggr].$ (13)
@@
$P\big(\hat{T}=t\big)=\penalty 10000\ C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB}e^{\mathbb{W}_{\!B}t}W_{BA}|x_{A}\rangle.$ (14)
This is a minor generalization of known results on first passage times skinner2021 . The factor of $\exp{\{\mathbb{W}_{\!B}t\}}$ takes into account the variable time spend in region $B$, integrating this factor over time provides $M:=-\mathbb{W}_{B}^{-1}$ which appeared in Eq. (9).
@@ II.3 Counting observables
The main feature of the framework proposed here is the characterization of statistics of counting observables within excursions. We first provide the definition of those observables and then discuss what happens once we restrict those to excursions, providing a few examples of important quantities.
For a given Markov jump process, we define a random variable $\hat{N}_{\ell xy}$ that counts the number of times the transition $y\to_{\ell}x$ was observed in one excursion. A counting observable is then defined as a linear combination of these random variables:
@@
with generic weights $\nu_{\ell xy}$ landi2024a . The counting observable registers the first transition of the excursion $A\to B$, all transitions within $B$, and the very last transition $B\to A$. Counting observables are defined irrespective of which state the excursion begins and ends.
With different choices of weights $\nu_{\ell xy}$ we can assess a wide variety of physical quantities. A typical quantity of interest is the number of times a specific transition $z^{\prime}\to z$ inside region $B$ occurs. This is obtained by setting $\nu_{\ell zz^{\prime}}=1$ and all other weights as zero. We provide analytical formulas for this case in Sec. IV.2. Similarly, $\nu_{\ell xy}=1\,\forall x,y$ for some $\ell$, can address the number of transitions of a given type. This can evaluate, for example, the activity of some hot reservoir in contrast to some other cold reservoir.
Another simple example of a counting observable is counting how many transitions occur in total. This is achieved by setting $\nu_{\ell xy}=1$ in Eq. (15). This counting observable is referred as the excursion activity $\hat{\mathcal{A}}$, it is related to the more usual dynamical activity (freneticity) which characterizes the number of transitions per unit time. The difference is that instead of characterizing activity per unit time, excursion activity accounts for activity per excursion. We provide analytical formulas for the excursion activity in Sec. IV.1. The dynamical activity has attracted a lot of interest because it plays a key role in uncertainty relations Di_Terlizzi2018 ; Nishiyama2024 ; Vo2022 . The simple relation between dynamical and excursion activity could be an important tool to explore excursions in the context of uncertainty relations.
@@
where $E_{x/y}$ are the energies of levels $x/y$ and $\beta_{\ell}$ is the inverse temperature of the reservoir associated with transition $\ell$. The counting observable representing the entropy produced in the process, given by the log of Eq. (16), can then be built with weights $\nu_{\ell xy}=-\beta_{\ell}(E_{x}-E_{y})$, which results in
@@
Other thermodynamic quantities can be constructed similarly. In Sec. V, we show that the entropy production satisfies a fluctuation theorem at the excursion level, where negative entropy production is exponentially suppressed, and use this result to derive a modified version of the thermodynamic uncertainty relation of Ref. Hasegawa2019 .
@@ III Statistics of counting observables within an excursion
$(\mathbb{W}_{\bm{\xi}})_{xy}=\begin{cases}\sum_{\ell}W_{\ell xy}e^{-i\sum_{\alpha}\nu_{\ell xy}^{\alpha}\xi_{\alpha}}&x\neq y\\[5.69046pt] -\Gamma_{x}&x=y,\end{cases}$ (18)
with $r$ counting fields $\xi_{\alpha}$. For an excursion starting in $x_{A}$ and ending in $y_{A}$, the joint probability that each observable $\hat{Q}_{\alpha}$ takes on a value $q_{\alpha}$ and that total excursion time is $\hat{T}=t$ reads
$P(\bm{q},t)=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi_{1}\ldots d\xi_{r}}{(2\pi)^{r}}\langle y_{A}|W_{AB\bm{\xi}}e^{\mathbb{W}_{\!B\bm{\xi}}t}W_{BA\bm{\xi}}|x_{A}\rangle e^{i\bm{q}\cdot\bm{\xi}},$ (19)
@@
$P(\bm{q})=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi_{1}\ldots d\xi_{r}}{(2\pi)^{r}}\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}W_{BA\bm{\xi}}|x_{A}\rangle e^{i\bm{q}\cdot\bm{\xi}},$ (20)
@@
[…]